glupin slWs a linear proportional relationship because it is a line that tontains the origin. O 20 40 60 80 100 Amount spent (S) YOUR TURN Determine if each of the following graphs represents a proportional or nonproportional relationship What do the slope and the y-intercept of the graph represent in this situation? 1. y 2. 4y 20 16 12 Persona Math Trai Online Assess 0 2 4 6 8 10 O 8 16 24 32 40 and Interves Lesson 4.4

these questions are based on proportional and nonportional situations using a graph. Please help.
Answer #1

Ist Graph:

Lets take the (x,y) pairs from the graph

x =1 ; y = 2 ; y/x = 2

x = 2 ; y = 3 ; y/x = 3/2 = 1.5

x = 3 ; y = 4 ; y/x = 4/3 = 1.33

All of the values of the ratios must be the same for the relationships to be proportional

So, it is non- proptional relationship.

IInd Graph :

x = 8 ; y = 2 ; y/x = 2/8 = 1/4

x= 16 ; y = 4 ; y/x = 4/16 = 1/4

x= 24; y = 6 ; y/x = 6/24 = 1/4

x =32 ; y = 8 ; y/x = 1/4

x = 40 ; y = 10 ; y/x = 1/4

All of the values of the ratios must be the same for the relationships to be proportional

Here the equivalent ration is1/4 for all values of x and y

So, relationship proptional